What is the 99th term of fibonacci sequence?


1
Expert's answer
2021-09-27T16:35:23-0400

Use Binet formula


"F_n=\dfrac{\alpha^n-\beta^n}{\alpha-\beta}"

"\alpha=\dfrac{1+\sqrt{5}}{2}, \beta=\dfrac{1-\sqrt{5}}{2}"

"\alpha-\beta=\sqrt{5}"

"F_{99}=\dfrac{(1+\sqrt{5})^{99}-(1-\sqrt{5})^{99}}{2^{99}\sqrt{5}}"

"F_{99}= 218922995834555169026"